OpenStudy (anonymous):

What mass of glucose (C6H12O6) is needed to prepare 200.0ml of a 2.50 M glucose solution

OpenStudy (anonymous):

Find the molar mass of glucose. Then\[\color{green}{\dfrac{2.50~\text{moles}}{1~\text{L}} = \dfrac{x~\text{moles}}{0.200~\text{L}}}\] Find the number of moles of glucose required (x). Finally \[\color{green}{\text{mass} = \text{number of moles} \times \text{molar mass}}\]